OpenStudy (anonymous):

MEDAL HELLP PLEASE What Is The Center and Radius of a circle (x-5)^2 /16 + (y-2)^2/16 = 1

OpenStudy (anonymous):

@bigballlz

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

this is of the form\[(x-a)^2+(y-b)^2=c^2\] for this circle center=(a,b) radius=c

OpenStudy (anonymous):

I know the radius is 4 but i have no clue what the center is

OpenStudy (sidsiddhartha):

\[\frac{ (x-5)^2 }{ 16 }+\frac{ (y-2)^2 }{ 16 }=1\]so it can be written as \[(x-5)^2+(y-2)^2=16\] now just comparing it with the above equation so center is=(5,2) and \[c^2=16\rightarrow c=4\] radius is 4

OpenStudy (anonymous):

THANKS Boo